OpenStudy (anonymous):

Use substitution to solve the system x = 2y + 3 and 4x - 5y = 9. What is the value of x? 2 1 -1 -2

OpenStudy (lilg132):

4(2y + 3) - 5y = 9 8y + 12 - 5y = 9 3y = 9 - 12 3y = -3 3y / 3 = -3 / 3 y = -1 4x - 5y = 9 4x - 5(-1) = 9 4x + 5 = 9 4x = 9 - 5 4x = 4 4x / 4 = 4 / 4 x = 1 remember to click good answer